How much energy is required to turn 27.03 g ice at 260k into water at 273k
1
Expert's answer
2018-10-08T07:51:08-0400

pecific heat of ice = 2.09 J/g·°C


Heat of fusion of water: ∆Hf = 334 J/g


1. Determine the heat required to raise the temperature of the ice from -13 °C (260 K) to 0 °C (273 K).


q1 = m*cice*∆T = 27.03g*2.09J/g·°C*(0°C – (-13°C)) = 734.41 J


2. Determine the heat required to convert 0 °C ice to 0 °C water.


q2 = m*∆Hf = 27.03g*334J/g = 9028.02 J


Therefore, the total heat required is


qtotal = q1 + q2 = 9762.43 J = 9.76 kJ